What is the final pH of a 100mL solution of 200mM histidine at pH of 6.0...

  1. What is the final pH of a 100mL solution of 200mM histidine at pH of 6.0 after adding 1L of 1mM HCl?  
Answer #1

The molar ratio between the salt and the acid components of the buffer is calculated, using the Henry Hasselbach equation cleared:

n Salt / n Acid = 10 ^ (pH - pKa) = 10 ^ (6 - 6.5) = 0.3

It has:

1) n Salt - 0.3 * n Acid = 0

2) n Salt + n Acid = M * V = 0.2 M * 0.1 L = 0.02 mol

System of equations between 1 and 2 is applied and you have:

n Salt = 0.005 mol

n Acid = 0.015 mol

The moles of HCl are calculated:

n HCl = 0.001 M * 1 L = 0.001 mol

The HCl reacts with the salt (decreasing it) and forms acid (increasing it), the pH of the new buffer is calculated:

pH = pKa + log (n Salt / n Acid) = 6.5 + log (0.005 - 0.001 / 0.015 + 0.001) = 5.9

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.